### v1

# Evaluating Scoring Bias in LLM-as-a-Judge

Abstract
The remarkable performance of Large Language Models (LLMs) gives rise to“LLM-as-a-Judge”, where LLMs are employed as evaluators for complex tasks. Moreover, it has been widely adopted across fields such as Natural Language Processing (NLP), preference learning, and various specific domains. However, there are various biases within LLM-as-a-Judge, which adversely affect the fairness and reliability of judgments. Current research on evaluating or mitigating bias in LLM-as-a-Judge predominantly focuses on comparison-based evaluations, while systematic investigations into bias in scoring-based evaluations remain limited. Therefore, we define scoring bias in LLM-as-a-Judge as the scores differ when scoring judge models are bias-related perturbed, and provide a well-designed framework to comprehensively evaluate scoring bias. We augment existing LLM-as-a-Judge benchmarks through data synthesis to construct our evaluation dataset and design multi-faceted evaluation metrics. Our experimental results demonstrate that the scoring stability of existing judge models is disrupted by scoring biases. Further exploratory experiments and discussions provide valuable insights into the design of scoring prompt templates and the mitigation of scoring biases on aspects such as score rubrics, score IDs, and reference answer selection. [^1] In the interest of protecting our intellectual property, specific technical details have been redacted from this preliminary version. The complete and unredacted manuscript is forthcoming.
Introduction
Large Language Models (LLMs), with outstanding performance, have been widely applied to Natural Language Generation (NLG) tasks such as summarization (Zhang et al. 2024) question answering (Robinson, Rytting, and Wingate 2022), and domain-specific dialogues (Wang et al. 2024a; Singhal et al. 2025). In these tasks, the model’s responses require find-grained evaluation of various aspects such as quality, factuality, and instruction-following, thereby providing valuable insights for model improvement. Traditional statistical NLG evaluation metrics, such as BLEU (Papineni et al. 2002) and ROUGE (Lin 2004), evaluate the generated text by n-gram overlap between the text and the golden answer. The above methods exhibit two primary limitations: (1) They focus solely on the similarity between the generated text and the golden answer, not quantifying intrinsic attributes of the text such as writing style, factuality, and educational value. (2) The resulting metrics lack interpretability, making it difficult to provide guidance for the model upgrade. There is an urgent need for fine-grained, multifaceted, and interpretable evaluation metrics and methods tailored to NLG tasks and LLM applications. With remarkable capabilities in instruction understanding and response generation, LLMs can provide diverse forms of evaluation output, such as scoring, comparison, and ranking. In addition, they can conduct evaluations along specified dimensions and generate critics for judgments, thus enhancing the interpretability of the evaluation. This line of research, referred to as “LLM-as-a-Judge” (Gu et al. 2024; Li et al. 2024a), has recently emerged as a prominent focus.
Figure 1: Our evaluation framework to measure the scoring bias in LLM-as-a-Judge.
Robustness is a critical factor in ensuring the fairness and reliability of LLM judgments, which requires that, under identical judge rubrics, LLM judges should provide consistent judgments for the same evaluation target. However, recent work (Park et al. 2024) indicates that LLM judges are subject to various biases, which undermine the self-consistency of their judgments. Many studies have sought to identify the types of bias that shift in judgments. (Wang et al. 2024b) finds that the judgments of the candidate responses can be easily hacked by altering their position in the context. (Wataoka, Takahashi, and Ri 2024) measures the self-preference bias in LLM-as-a-Judge, which refers to the tendency of LLM judges to overestimate the quality of their own outputs. (Wei et al. 2024) considers length bias as LLM judges favor longer responses or shorter responses.
Although existing studies systematically investigated various biases in LLM-as-a-Judge, several pressing issues remain to be addressed. Firstly, while some work (Wang et al. 2024b; Ye et al. 2024a) has addressed biases such as positional bias and refinement awareness in scoring-based LLM judges, many existing research remains focused on biases in comparative LLM judges. Existing work (Ye et al. 2024a) attempts to investigate specific biases in scoring tasks, such as self-enhancement bias and refinement-aware bias. However, there is still a lack of comprehensive research on biases in scoring-based LLM-as-a-Judge. Secondly, the biases defined in existing work are predominantly related to the attributes or presentation forms of the evaluation targets, such as position, length, and response model. However, variation inherent in the prompt of LLM-based judge itself may also impact the self-consistency of judgments, which lacks a systematic investigation in current research.
To solve the above problems, we propose scoring bias, defined as the scoring shifts of LLM-as-a-Judge under perturbations of the scoring prompt. In this work, we propose three types of scoring bias: score rubrics order bias, score IDs bias, and reference answer score bias, and design a framework to systematically evaluate those biases. We select four existing LLM-as-a-Judge benchmarks as evaluation datasets and develop a series of metrics to quantify the robustness of LLM judges against the score bias. Our key contributions can be summarized as follows:
  • We pioneer a systematic evaluation of bias in scoring-based LLM-as-a-Judge with definition of three types of scoring bias and five perturbation methods.
  • Through automatic data synthesis, we expand existing LLM-as-a-Judge benchmarks into our evaluation dataset without human labor costs. Furthermore, we evaluate scoring bias from multiple perspectives through a comprehensive design of evaluation metrics.
  • Our comprehensive experiments and in-depth analysis find that scoring bias affects the robustness and reliability of various existing judge models. We also provide valuable recommendations for the design of scoring prompts and the mitigation of scoring bias.
Related Work
LLM-as-a-Judge
(Gu et al. 2024) defines LLM-as-a-Judge as the task in which an LLM generates an evaluation result in any form given input data and context. Common forms of LLM judgment include score generation (Gao et al. 2023), pairwise comparison (Qin et al. 2024), ranking (Li, Patel, and Du 2023), multiple-choice selection (Li et al. 2024b), and critic generation (Ke et al. 2024). Improvement strategies for LLM-as-a-Judge include meticulous design of evaluation prompts (Liu et al. 2023), improvement of LLM’s abilities such as fine-tuning (Wang et al. 2024c), multi-LLM collaboration pipelines such as feedback mechanism (Xu et al. 2023), and output optimization strategies such as multi-output ensemble (Sottana et al. 2023). The performance of LLM-as-a-Judge is usually measured using metrics that represent its alignment with human annotators or advanced LLM judgments such as agreement (Thakur et al. 2024), Cohen’s Kappa, and Spearman’s correlation (Bai et al. 2023). LLM-as-a-Judge is widely applied to evaluation (Lin and Chen 2023), alignment (Bai et al. 2022), retrieval (Sun et al. 2023), reasoning (Gao et al. 2024a), and other specific domains (Ryu et al. 2023; Babaei and Giudici 2024). LLM-as-a-Judge faces challenges in bias and vulnerability, which significantly compromise the fairness and reliability of LLM judges. (Li et al. 2024a)
Bias in LLM-as-a-Judge
(Ye et al. 2024a) defines the prompt for LLM judges as three components: system instruction, question, and response. Based on this, they define bias in LLM-as-a-Judge as that the judgments differ when system instruction or responses is bias-related modified. Common types of bias in LLM-as-a-Judge include position bias, length bias, self-enhancement bias, etc. Those biases are often incorporated into the evaluation scope of existing LLM-as-a-Judge benchmarks. (Zheng et al. 2023; Wei et al. 2024) Currently, the types of bias still remain a focal point of investigation, and numerous benchmarks are proposed to evaluate bias in LLM-as-a-Judge. OffsetBias (Park et al. 2024) points out six types of bias in LLM-as-a-Judge including length bias, concreteness bias, empty reference bias, and so on. CALM (Ye et al. 2024a) quantifies and analyzes 12 types of bias in LLM judges. However, the biases defined in current work are predominantly focused on comparative evaluation tasks, with a lack of systematic research and definitions specifically addressing biases in scoring-based evaluation tasks. Similar to research on adversarial attacks, the evaluation metrics for bias in LLM-as-a-Judge are often based on the consistency between bias-related modified judgments and the original judgments, or on alignment with golden answers. Some studies focus on mitigating bias in LLM-as-a-Judge and aim to develop more robust LLM judges. (Liu et al. 2024c) introduces Pairwise-Preference Search (PAIRS) to mitigate biases in LLM evaluators. (Wang et al. 2024b) proposes three methods to calibrate positional bias in LLM-as-a-Judge, including multiple evidence calibration, balanced position calibration, and human-in-the-loop calibration. Our evaluation results on scoring bias will serve as a reference to guide future research in mitigating scoring biases in judge models and enhancing the robustness of scoring-based evaluations.
Evaluation Method
Figure 2: The scoring prompt template with four components from (Kim et al. 2024c). The red text represents the components may be perturbed.
Figure 1 shows our evaluation framework, inspired by (Ye et al. 2024a).
Dataset Data Size Human Judgment Reference Answer
BiGGen Bench 2,780
FLASK 2,001
MT Bench 320
Vicuna Bench 320
Table 1: Statics of our evaluation datasets.
We evaluate the scoring bias in LLM-as-a-Judge by perturbing the scoring prompt and observing whether the scoring judgments are affected, which follows (Kim et al. 2024c).
Bias Types
We define three types of bias targeting the key components related to scoring in LLM judges:
  • Score Rubrics Order
  • Score IDs
  • Reference Answer Score
Datasets
Following previous work (Kim et al. 2024c; Chiang, Lee, and Lukasik 2025), we select four LLM-as-a-Judge datasets for scoring bias evaluation:
  • BiGGen Bench (Kim et al. 2024b): This is a principled generation benchmark with instance-specific evaluation criteria. We use its test set for non-multilingual capabilities with 2,780 responses from 695 instances that are evaluated by human annotators.
  • FLASK (Ye et al. 2024b): This is a fine-grained evaluation benchmark with 200 test prompts, 12 score rubrics, and 2,001 responses. It also includes scores labeled by human annotators.
  • MT Bench (Zheng et al. 2023): This is a multi-turn chat benchmark with 80 test prompts. (Kim et al. 2024a) write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B (Xu et al. 2024), Vicuna-13B (Chiang et al. 2023), Llama-2-Chat-13B (Touvron et al. 2023), and GPT-3.5-Turbo-0613 (OpenAI 2023a).
  • Vicuna Bench (Chiang et al. 2023): This is a single-turn chat benchmark with 80 test prompts.(Kim et al. 2024a) also write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B, Vicuna-13B, Llama-2-Chat-13B, and GPT-3.5-Turbo-0613.
Model BiGGen Bench FLASK
$\rho$ $r$ $\rho$ $r$
GPT-4.1 0.6048 0.6360 0.6401 0.6823
GPT-4o 0.6047 0.6340 0.6340 0.6557
GPT-4 0.5741 0.6045 0.6207 0.6862
Table 2: Correlations between LLM judgement pipeline and human when generating golden scores. $\rho$ is Spearman’s rank correlation coefficient. $r$ is Pearson’s correlation coefficient. The GPT-4.1 and GPT-4o scores are generated by our work, the GPT-4 scores are derived from the original datasets.
The detailed statics of our selected datasets are shown in Table 1. For advanced LLM scores, the above datasets have evaluation results obtained from GPT-4 (OpenAI 2023b). Due to the earlier release of GPT-4, we use more recent models GPT-4.1 (OpenAI 2025) and GPT-4o (OpenAI 2024) to obtain golden scores. Inspired by (Kim et al. 2024a, c), we repeat the scoring three times through a judge prompt with a reference answer that marked a score of 5 and make a majority vote to obtain the final score from LLM evaluators. Table 2 shows the correlation between advanced LLM scores and human scores. The results show that the GPT-4.1 scores exhibit the highest correlation with human scores, so we select the GPT-4.1 scores as the golden LLM scores.
Each instance in BiGGen Bench and FLASK is accompanied by a reference answer with score 5. To measure the impact of the reference answer score on the judgment robustness, it is necessary to additionally generate reference answers with scores ranging from 1 to 4. We propose a generation-review pipeline to synthesize reference answers for each sample in FLASK, as shown in Figure 3. Given an instruction and score rubrics, the response model generates the response according to the description from score 1 to 4. Then a review model checks whether the response matches the corresponding score description. If match, we accept the generated responses; otherwise, we rerun the pipeline until a match is achieved. We alternately apply GPT-4.1 and GPT-4o as response model and review model in the pipeline.
Metrics
Figure 3: The proposed generation-review pipeline for synthesizing reference answers.
To evaluate the scoring bias in LLM-as-a-Judge, we consider the impact or perturbation methods on scoring accuracy and scoring tendency.
Experiments
Model Bias Perturbation BiGGen Bench FLASK MT Bench Vicuna Bench
Human GPT-4.1 Human GPT-4.1 GPT-4.1 GPT-4.1
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
GPT-4o - - 0.5685 0.6165 0.7572 0.8059 0.5529 0.5914 0.7369 0.7676 0.7578 0.7724 0.6862 0.7696
Rubrics Order Descending 0.5852 0.6245 0.7787 0.8233 0.5754 0.6080 0.7602 0.7760 0.7479 0.7597 0.6734 0.7635
Random 0.5963 0.6313 0.7698 0.8124 0.5592 0.5893 0.7420 0.7573 0.7673 0.7838 0.6701 0.7504
Score IDs Letter 0.5641 0.6034 0.7595 0.7994 0.5558 0.5977 0.7341 0.7609 0.7317 0.7454 0.6539 0.7712
Roman numerals 0.5754 0.6171 0.7709 0.8148 0.5535 0.5912 0.7380 0.7641 0.7725 0.7786 0.7073 0.7821
DeepSeek-V3 - - 0.5314 0.5748 0.6775 0.7394 0.5163 0.5660 0.6711 0.7063 0.6323 0.6593 0.4313 0.5443
Rubrics Order Descending 0.5356 0.5812 0.6611 0.7211 0.5181 0.5554 0.6555 0.6858 0.6183 0.6436 0.4993 0.5854
Random 0.5136 0.5609 0.6547 0.7125 0.4999 0.5470 0.6521 0.6752 0.6178 0.6460 0.5086 0.5482
Score IDs Letter 0.5378 0.5689 0.6803 0.7179 0.5314 0.5755 0.6663 0.6999 0.6489 0.6715 0.4832 0.5657
Roman numerals 0.5164 0.5636 0.6688 0.7195 0.5066 0.5542 0.6750 0.7109 0.6558 0.6642 0.4416 0.5285
Qwen3-32B - - 0.5421 0.5855 0.7328 0.7715 0.5323 0.5695 0.6549 0.6807 0.6959 0.7053 0.6419 0.7188
Rubrics Order Descending 0.5572 0.5989 0.7443 0.7862 0.5345 0.5657 0.6586 0.6839 0.6843 0.6989 0.6205 0.7020
Random 0.5528 0.5938 0.7404 0.7824 0.5115 0.5456 0.6537 0.6748 0.6554 0.6688 0.6416 0.7252
Score IDs Letter 0.5353 0.5851 0.7280 0.7770 0.5393 0.5724 0.6650 0.6975 0.6795 0.6930 0.6157 0.6808
Roman numerals 0.5139 0.5629 0.6992 0.7412 0.5257 0.5781 0.6630 0.6980 0.6360 0.6418 0.6072 0.7126
Qwen3-8B - - 0.5092 0.5308 0.6522 0.6492 0.5035 0.5309 0.6464 0.6616 0.6238 0.6342 0.6639 0.7284
Rubrics Order Descending 0.5241 0.5456 0.6458 0.6559 0.4885 0.5339 0.6315 0.6566 0.6005 0.6135 0.5500 0.6647
Random 0.4869 0.5025 0.6091 0.5950 0.4778 0.5004 0.6061 0.6216 0.5949 0.5983 0.5135 0.6053
Score IDs Letter 0.4853 0.5117 0.6410 0.6665 0.4873 0.5244 0.6334 0.6586 0.5985 0.6126 0.5812 0.6890
Roman numerals 0.4801 0.5120 0.6231 0.6445 0.4606 0.5087 0.6177 0.6468 0.6016 0.5975 0.6111 0.7178
Mistral-24B - - 0.5059 0.5495 0.6910 0.7201 0.4950 0.5385 0.6303 0.6593 0.5943 0.5762 0.6116 0.6917
Rubrics Order Descending 0.5369 0.5761 0.6987 0.7213 0.4922 0.5327 0.6362 0.6534 0.6104 0.6102 0.6137 0.7082
Random 0.5053 0.5450 0.6569 0.6841 0.4756 0.5042 0.6212 0.6338 0.5603 0.5602 0.5461 0.6418
Score IDs Letter 0.5114 0.5632 0.6685 0.7084 0.4880 0.5339 0.6311 0.6594 0.5375 0.5485 0.5558 0.6608
Roman numerals 0.5102 0.5590 0.6640 0.6962 0.4957 0.5494 0.6533 0.6872 0.5863 0.5907 0.5903 0.6905
Table 3: Correlation between golden scores and judge model scores with bias in score rubrics order and score IDs. For each judge model, the first row is the metrics without bias, the remaining rows is the metrics under different types of bias. The differences between results with and without bias reflect the extent to which bias impacts the scoring robustness. DeepSeek-V3 is DeepSeek-V3-671B. Mistral-24B is Mistral-Small-24B-Instruct. $\rho$ is Spearman’s rank correlation coefficient. $r$ is Pearson’s correlation coefficient.
Setup
Judge Models.
Previous work (Gao et al. 2024b; Li et al. 2024c; Liu et al. 2024b) has reached a consensus that judge models should be stronger models to ensure evaluation reliability. (Ye et al. 2024a) However, as an exploratory study, we think that capability and size of the judge model should also be considered as influence factors. Therefore, we select diverse judge models that include closed-source, open-source, large-size, and small-size models. They are GPT-4o, DeepSeek-V3-671B (Liu et al. 2024a), Qwen3-32B, Qwen3-8B (Yang et al. 2025), and Mistral-Small-24B-Instruct-2501 (MistralAI 2025). Those judge models we selected differ from the models used to generate response in the evaluation datasets, thus mitigating potential self-enhancement bias (Zheng et al. 2023; Ye et al. 2024a).
Judge Prompt.
The prompt we use for scoring-based judgment is shown in Figure 2, following (Kim et al. 2024c). The unperturbed prompt is the one with ascending score rubrics, Arabic numeral score IDs, and no reference answer.
Hyperparameters.
As our goals for scoring bias evaluations, we aim to ensure the stability of the model’s generative capabilities. Therefore, we set the judge models’ temperature to 0 while scoring.
Main Results
Judge models are sensitive to scoring biases in input prompt templates.
Table 3 and Figure 4 show the correlation of judge model scores and golden scores on all perturbation methods. All judge models we evaluated, even for an advanced model such as GPT-4, fluctuations in score results are observed under different biases, indicating that scoring bias has a significant impact on the robustness of LLM-as-a-Judge. Furthermore, the impact exhibits certain differences depending on the choice of judge model:
(1) The capability and size of the judge models are critical factors influencing the robustness against scoring bias. Consistent with existing consensus (Gao et al. 2024b; Li et al. 2024c; Liu et al. 2024b), our experimental results indicate that judge models with advanced capability and larger size tend to exhibit higher reliability on scoring. As shown in Table 3, for example, across all perturbation methods and datasets, the fluctuation in correlation coefficients between GPT-4o and the golden scores remains with in 0.03, except in introduction of a reference answer scored 5. By comparing the results of Qwen3-32B and Qwen3-8B, it is evident that judge models with larger parameter sizes demonstrate greater robustness in scoring. In comparison with the GPT-4.1 scores in BiGGen-Bench and Vicuna Bench, the correlation coefficients of the Qwen3-8B scores exhibit maximum fluctuations exceeding 0.05 under score rubrics order bias and score IDs bias. With reference answer score bias, the maximum fluctuation reaches nearly 0.2. However, the changes in the correlation coefficient of Qwen3-32B under biases are generally within 0.03. Therefore, more capable judge models should be prioritized in scoring evaluations.
Figure 4: Spearman’s correlation between human annotator scores and judge model scores with bias in reference answer score on FLASK.
(2) Different judge models exhibit varying sensitivity to scoring bias. The impact of bias on the scoring accuracy of judge models can be either positive or negative. However, this impact is not fixed; different judge models have different trends in scoring accuracy when affected by the same bias. As Table 3 shows, for example, in most cases, the correlation between GPT-4o scores and golden scores increases when Roman numerals are used as scoring IDs. But when DeepSeek-V3-671B, Qwen3-32B and Qwen3-8B are employed as judge models, the use of Roman numerals generally exerts a negative effect on their scoring accuracy. An interesting phenomenon is that the impact of reference answer score bias on the judge models we evaluated appears to be consistent, which will be further analyzed in subsequent sections. Our experimental results can serve as a valuable reference for the selection of judge models and the design of scoring prompt templates.
(3) Judge models have their own scoring tendencies. Figure 5 presents the scoring tendencies of different judge models under various scoring biases on the same dataset FLASK in the form of heatmaps. It is obvious that each judge model exhibits its own specific scoring preferences. Under unbiased conditions, DeepSeek-V3-671B assigns a score of 5 to more than half of the instruction-response pairs. Compared to the human scoring distribution, Mistral-Small-24B-Instruct has a stronger tendency to score a 4 to the samples. Other judge models show smoother scoring distribution shifts, but certain scores are still preferred. GPT-4o demonstrates a preference for score 4, while Qwen3-32B favors score 5. An examination of Figure 5 reveals that score rubrics order bias and score IDs bias have a certain impact on the scoring tendencies of the judge models; however, they do not fundamentally alter the overall score distribution. An exception is Qwen3-8B, whose score distribution is more significantly influenced by biases, possibly due to its small size. It is particularly evident that the reference answer score bias exerts a substantial impact on the score distribution, which will be analyzed in detail in the following sections. Our experimental results show that judge models inherently exhibit biases in their scoring tendencies. Therefore, debiasing the judge models is crucial to ensuring the fairness and reliability of the evaluation.
Figure 5: Heatmaps of judge models’ score distribution with all bias types on FLASK. The horizontal axis is the score, where -1 represents the failed judgment. The vertical axis is the perturbation method. The top lines of all heatmaps are the human scoring tendencies for comparison, and the second lines are results from unbiased judge models.
Exploratory Experiments and Discussion
For advanced or large size judge models, application of scoring prompt that differ from usual preferences are worth to be considered.
In almost all scoring-based evaluation tasks, Arabic numerals are commonly used as scoring IDs. The score rubrics order typically follows a sequential order from score 1 to score 5. These choices align with human writing conventions and cognitive preferences in the design of scoring prompts. However, our experimental results shown in Table 3 indicate that the scoring prompts designed to align with the usual preferences do not necessarily produce the most accurate scoring results. As mentioned earlier, the performance of the GPT-4o judge model in scoring evaluations improves when Roman numerals are used as scoring IDs. In more than half of the cases, using letter grades as scoring Ids improves the scoring accuracy of DeepSeek-V3-671B. Regarding the score rubrics order bias, adopting a descending order improves the scoring accuracy of GPT-4o, Qwen3-32B, and Mistral-Small-24B-Instruct in most instances. Intuitively, employing a random scoring rubrics order often tends to negatively impact the scoring accuracy of all judge models. Based on the above analysis, we recommend incorporating certain intentional deviations from conventional practices in the design of scoring prompts.
The reference answer score bias perform intriguing impacts on the robustness of scoring-based evaluation and play a critical role.
The impact of reference answer score bias on the scoring accuracy and scoring tendency of the judge model is presented in Figure 4 and Figure 5. The experimental results indicate that, compared to the score rubrics order bias and the score IDs bias, the influence of the reference answer score bias on scoring robustness is both more pronounced and demonstrates stronger regularity. Figure 4 illustrates the impact of introducing reference answers with different scores in the input prompt on the scoring accuracy of the judge model. For GPT-4o, DeepSeek-V3-671B, and Mistral-Small-24B-Instruct, the scoring accuracy in most cases improves when reference answers are introduced. However, for the two Qwen judge models, only reference answers with scores of 4 or 5 contribute to enhancing scoring performance. Figure 4 also reveals an interesting phenomenon: the higher the score of the reference answer, the better the scoring accuracy of the judge model. Although DeepSeek-V3-671B, Qwen3-32B, and Mistral-Small-24B-Instruct do not strictly adhere to this pattern, the correlation between their scores and human scores similarly exhibits an upward trend as the reference answer score increases.
Figure 5 shows that the score of the reference answer has a significant impact on the scoring tendency of the judge model. Specifically, the scoring results from DeepSeek-V3-671B, Qwen3-32B, and Qwen3-8B appear to be highly influenced by the score of the reference answer. When the score of the reference answer falls within the range of 1 to 4, these models tend to assign similar scores to the evaluated instruction-response pairs. Interestingly, when a reference answer with a score of 5 is introduced, the scoring behavior of the judge model appears to become more rational. This is particularly notable for DeepSeek-V3-671B, which exhibits a strong tendency to assign a score of 5 without bias. However, the introduction of a reference answer scored at 5 mitigates this phenomenon.
In summary, the above analysis reveals that the score of the reference answer introduced in the prompt has a significant impact on the judge model’s scoring behavior. In particular, when reference answers with a score of 5 are introduced, the models achieve the highest scoring accuracy without disproportionately influencing their scoring preferences. Therefore, it can be concluded that when incorporating reference answers into the scoring-based judge prompt, a full-marked reference answer is the best choice. Interestingly, this conceptually aligns with the principles of traditional statistical metrics, such as BLEU and ROUGE, which emphasize the critical role of golden answers in evaluation.
Limitations and Future Work
Although our work offers notable conclusions and insights, some limitations still remain for future work.
Mitigating scoring bias needs further exploration. Although there are no dedicated and fancy methods for addressing scoring bias, intuitive and straightforward approaches, such as scoring multiple times followed by majority voting or averaging the scores, require empirical validation for their effectiveness in mitigating scoring bias.
Additionally, although our work demonstrates the existence of scoring bias in LLM-as-a-Judge, the underlying causes of scoring bias remain to be validated. Approaches such as training data analysis and information flow observation (Yuan et al. 2025) may help identify the reason scoring bias exists within or outside the model.
Conclusion
In this work, we propose scoring bias in LLM-as-a-Judge, defined as scoring judgments shift when judge models are perturbed. We design an evaluation framework, including three types of scoring bias: score rubrics order bias, score IDs bias, and reference answer score bias. We select four LLM-as-a-Judge benchmarks and expand them as our evaluation datasets through data synthesis. Our evaluation metrics can quantify the impact of scoring bias on the scoring accuracy and scoring tendencies of the judge models. Our experimental results clearly demonstrate that scoring bias can affect the robustness and reliability of existing scoring-based judge models. Judge models with different capabilities, sizes, and architectures exhibit different sensitivities to various scoring biases. Our further exploratory experiments and discussions reveal several intriguing phenomena. In summary, our work provides detailed evidence and valuable recommendations for evaluating and improving the scoring reliability of LLM-as-a-Judge.

### v2

@@
 Abstract
-The remarkable performance of Large Language Models (LLMs) gives rise to“LLM-as-a-Judge”, where LLMs are employed as evaluators for complex tasks.
-Moreover, it has been widely adopted across fields such as Natural Language Processing (NLP), preference learning, and various specific domains.
-However, there are various biases within LLM-as-a-Judge, which adversely affect the fairness and reliability of judgments.
-Current research on evaluating or mitigating bias in LLM-as-a-Judge predominantly focuses on comparison-based evaluations, while systematic investigations into bias in scoring-based evaluations remain limited.
-Therefore, we define scoring bias in LLM-as-a-Judge as the scores differ when scoring judge models are bias-related perturbed, and provide a well-designed framework to comprehensively evaluate scoring bias.
-We augment existing LLM-as-a-Judge benchmarks through data synthesis to construct our evaluation dataset and design multi-faceted evaluation metrics.
-Our experimental results demonstrate that the scoring stability of existing judge models is disrupted by scoring biases.
-Further exploratory experiments and discussions provide valuable insights into the design of scoring prompt templates and the mitigation of scoring biases on aspects such as score rubrics, score IDs, and reference answer selection.
-[^1] In the interest of protecting our intellectual property, specific technical details have been redacted from this preliminary version.
-The complete and unredacted manuscript is forthcoming.
+The remarkable capabilities of Large Language Models (LLMs) have led to the emergence of “LLM-as-a-Judge”, a paradigm where LLMs serve as automated evaluators for complex tasks.
+This approach has been widely adopted across diverse fields, including natural language processing and preference learning.
+However, the fairness and reliability of these LLM judges are often compromised by inherent biases.
+While current research has predominantly focused on biases in comparison-based evaluations, systematic investigations into scoring-based evaluations remain limited.
+To address this gap, we first formally define scoring bias as the discrepancy in scores when the judge model’s prompt is exposed to bias-inducing perturbations.
+We then propose a comprehensive framework to evaluate this bias.
+This framework is supported by two key contributions: (1) new evaluation datasets constructed by augmenting existing benchmarks with synthesized data, and (2) a suite of multi-faceted metrics designed for scoring bias analysis.
+Our experiments reveal that the scoring stability of LLM judges is significantly disrupted by these biases.
+Furthermore, our findings offer actionable insights for designing more robust scoring prompts and mitigating biases related to score rubrics, score IDs, and reference answer selection.
 Introduction
 Large Language Models (LLMs), with outstanding performance, have been widely applied to Natural Language Generation (NLG) tasks such as summarization (Zhang et al.
-2024) question answering (Robinson, Rytting, and Wingate 2022), and domain-specific dialogues (Wang et al.
+2024), question answering (Robinson, Rytting, and Wingate 2022), and domain-specific dialogues (Wang et al.
 2024a; Singhal et al.
 2025).
-In these tasks, the model’s responses require find-grained evaluation of various aspects such as quality, factuality, and instruction-following, thereby providing valuable insights for model improvement.
+In these tasks, the model’s responses require fine-grained evaluation of various aspects such as quality, factuality, and instruction-following, thereby providing valuable insights for model improvement.
 Traditional statistical NLG evaluation metrics, such as BLEU (Papineni et al.
-2002) and ROUGE (Lin 2004), evaluate the generated text by n-gram overlap between the text and the golden answer.
-The above methods exhibit two primary limitations: (1) They focus solely on the similarity between the generated text and the golden answer, not quantifying intrinsic attributes of the text such as writing style, factuality, and educational value.
-(2) The resulting metrics lack interpretability, making it difficult to provide guidance for the model upgrade.
-There is an urgent need for fine-grained, multifaceted, and interpretable evaluation metrics and methods tailored to NLG tasks and LLM applications.
+2002) and ROUGE (Lin 2004), evaluate the generated text by measuring n-gram overlap between the text and the golden answer.
+The above methods exhibit two primary limitations: (1) They focus solely on the similarity between the generated text and the golden answer, rather than quantifying intrinsic attributes of the text such as writing style, factuality, and educational value.
+(2) The resulting metrics lack interpretability, making it difficult to provide guidance for the model improvement.
+There is an urgent need for fine-grained, multi-faceted, and interpretable evaluation metrics and methods tailored to NLG tasks and LLM applications.
 With remarkable capabilities in instruction understanding and response generation, LLMs can provide diverse forms of evaluation output, such as scoring, comparison, and ranking.
-In addition, they can conduct evaluations along specified dimensions and generate critics for judgments, thus enhancing the interpretability of the evaluation.
+In addition, they can conduct evaluations along specified dimensions and generate critiques for their judgments, thus enhancing the interpretability of the evaluation.
 This line of research, referred to as “LLM-as-a-Judge” (Gu et al.
@@
 2024a), has recently emerged as a prominent focus.
+Figure 1: Even slight perturbations to LLM judges lead to inconsistent scores for the same response.
 Figure 2: Our evaluation framework to measure the scoring bias in LLM-as-a-Judge.
-Robustness is a critical factor in ensuring the fairness and reliability of LLM judgments, which requires that, under identical judge rubrics, LLM judges should provide consistent judgments for the same evaluation target.
+Robustness is a critical factor in ensuring the fairness and reliability of LLM judgments, which requires that, under identical judge rubrics, LLM judges should provide consistent judgments for semantically equivalent but differently phrased evaluation targets.
 However, recent work (Park et al.
 2024) indicates that LLM judges are subject to various biases, which undermine the self-consistency of their judgments.
-Many studies have sought to identify the types of bias that shift in judgments.
-(Wang et al.
-2024b) finds that the judgments of the candidate responses can be easily hacked by altering their position in the context.
-(Wataoka, Takahashi, and Ri 2024) measures the self-preference bias in LLM-as-a-Judge, which refers to the tendency of LLM judges to overestimate the quality of their own outputs.
-(Wei et al.
-2024) considers length bias as LLM judges favor longer responses or shorter responses.
-Although existing studies systematically investigated various biases in LLM-as-a-Judge, several pressing issues remain to be addressed.
+Many studies have sought to identify the types of bias that cause shifts in judgments.
+Wang et al.
+finds that the judgments of the candidate responses can be easily influenced by altering their position in the context.
+Wataoka, Takahashi, and Ri measures the self-preference bias in LLM-as-a-Judge, which refers to the tendency of LLM judges to overestimate the quality of their own outputs.
+Wei et al.
+defines length bias as the tendency of LLM judges to favor longer or shorter responses.
+Although existing studies have systematically investigated various biases in LLM-as-a-Judge, several pressing issues remain to be addressed.
 Firstly, while some work (Wang et al.
 2024b; Ye et al.
-2024a) has addressed biases such as positional bias and refinement awareness in scoring-based LLM judges, many existing research remains focused on biases in comparative LLM judges.
+2024a) has addressed biases such as positional bias and refinement awareness in scoring-based LLM judges, many existing studies remain focused on biases in comparative LLM judges.
 Existing work (Ye et al.
@@
 However, there is still a lack of comprehensive research on biases in scoring-based LLM-as-a-Judge.
-Secondly, the biases defined in existing work are predominantly related to the attributes or presentation forms of the evaluation targets, such as position, length, and response model.
-However, variation inherent in the prompt of LLM-based judge itself may also impact the self-consistency of judgments, which lacks a systematic investigation in current research.
-To solve the above problems, we propose scoring bias, defined as the scoring shifts of LLM-as-a-Judge under perturbations of the scoring prompt.
-In this work, we propose three types of scoring bias: score rubrics order bias, score IDs bias, and reference answer score bias, and design a framework to systematically evaluate those biases.
-We select four existing LLM-as-a-Judge benchmarks as evaluation datasets and develop a series of metrics to quantify the robustness of LLM judges against the score bias.
+Secondly, the biases defined in existing work are predominantly related to the choice of response models and the attributes or presentation forms of the evaluation targets, such as position and length.
+However, variation inherent in the LLM judge itself, such as the ordering of score rubrics and the selection of reference answers, may also impact the self-consistency of judgments, which lacks a systematic investigation in current research.
+To solve the above problems, we propose scoring bias, defined as the scoring shifts of LLM-as-a-Judge under perturbations of the scoring prompt, as Figure 1 shows.
+In this work, we propose three types of scoring bias: score rubric order bias, score ID bias, and reference answer score bias, and design a framework to systematically evaluate these biases.
+We select four existing LLM-as-a-Judge benchmarks as evaluation datasets and utilize correlation with golden scores and score distributions as metrics to quantify the robustness and scoring tendency of LLM judges.
 Our key contributions can be summarized as follows:
-  • We pioneer a systematic evaluation of bias in scoring-based LLM-as-a-Judge with definition of three types of scoring bias and five perturbation methods.
-  • Through automatic data synthesis, we expand existing LLM-as-a-Judge benchmarks into our evaluation dataset without human labor costs.
-Furthermore, we evaluate scoring bias from multiple perspectives through a comprehensive design of evaluation metrics.
-  • Our comprehensive experiments and in-depth analysis find that scoring bias affects the robustness and reliability of various existing judge models.
-We also provide valuable recommendations for the design of scoring prompts and the mitigation of scoring bias.
+  • We conduct the first systematic evaluation of scoring bias in LLM-as-a-Judge, introducing a formal definition of this bias and a comprehensive methodology to quantify it from multiple perspectives.
+  • To address the critical lack of datasets for scoring bias analysis, we introduce an automatic data synthesis pipeline that expands existing LLM-as-a-Judge benchmarks into a tailored evaluation corpus, thereby establishing a foundational resource for future studies.
+  • We empirically demonstrate that even the most advanced LLMs suffer from substantial scoring biases and exhibit obvious self-bias tendency.
+Furthermore, our analysis yields valuable recommendations for designing more robust scoring prompts and mitigating scoring biases.
 Related Work
 LLM-as-a-Judge
-(Gu et al.
-2024) defines LLM-as-a-Judge as the task in which an LLM generates an evaluation result in any form given input data and context.
+Gu et al.
+defines LLM-as-a-Judge as the task in which an LLM generates an evaluation result in any form given input data and context.
 Common forms of LLM judgment include score generation (Gao et al.
@@
 2024), ranking (Li, Patel, and Du 2023), multiple-choice selection (Li et al.
-2024b), and critic generation (Ke et al.
+2024b), and critique generation (Ke et al.
 2024).
@@
 2023), improvement of LLM’s abilities such as fine-tuning (Wang et al.
-2024c), multi-LLM collaboration pipelines such as feedback mechanism (Xu et al.
-2023), and output optimization strategies such as multi-output ensemble (Sottana et al.
+2024c), multi-LLM collaboration pipelines such as feedback mechanisms (Xu et al.
+2023), and output optimization strategies such as multi-output ensembling (Sottana et al.
 2023).
@@
 2023; Babaei and Giudici 2024).
-LLM-as-a-Judge faces challenges in bias and vulnerability, which significantly compromise the fairness and reliability of LLM judges.
-(Li et al.
-2024a)
+LLM-as-a-Judge faces challenges of bias and vulnerability, which significantly compromise the fairness and reliability of LLM judges (Li et al.
+2024a).
 Bias in LLM-as-a-Judge
-(Ye et al.
-2024a) defines the prompt for LLM judges as three components: system instruction, question, and response.
-Based on this, they define bias in LLM-as-a-Judge as that the judgments differ when system instruction or responses is bias-related modified.
+Ye et al.
+defines the prompt for LLM judges as three components: system instruction, question, and response.
+Based on this, they define bias in LLM-as-a-Judge as the tendency for judgments to differ when the system instruction or responses are modified in a bias-related manner.
 Common types of bias in LLM-as-a-Judge include position bias, length bias, self-enhancement bias, etc.
-Those biases are often incorporated into the evaluation scope of existing LLM-as-a-Judge benchmarks.
+These biases are often incorporated into the evaluation scope of existing LLM-as-a-Judge benchmarks.
 (Zheng et al.
 2023; Wei et al.
-2024) Currently, the types of bias still remain a focal point of investigation, and numerous benchmarks are proposed to evaluate bias in LLM-as-a-Judge.
+2024) Currently, the types of bias remain a focal point of investigation, and numerous benchmarks have been proposed to evaluate bias in LLM-as-a-Judge.
 OffsetBias (Park et al.
-2024) points out six types of bias in LLM-as-a-Judge including length bias, concreteness bias, empty reference bias, and so on.
+2024) identifies six types of bias in LLM-as-a-Judge including length bias, concreteness bias, empty reference bias, and so on.
 CALM (Ye et al.
 2024a) quantifies and analyzes 12 types of bias in LLM judges.
-However, the biases defined in current work are predominantly focused on comparative evaluation tasks, with a lack of systematic research and definitions specifically addressing biases in scoring-based evaluation tasks.
+However, the biases defined in current work are predominantly focused on comparative evaluation tasks, and there is a lack of systematic research and definitions specifically addressing biases in scoring-based evaluation tasks.
 Similar to research on adversarial attacks, the evaluation metrics for bias in LLM-as-a-Judge are often based on the consistency between bias-related modified judgments and the original judgments, or on alignment with golden answers.
 Some studies focus on mitigating bias in LLM-as-a-Judge and aim to develop more robust LLM judges.
-(Liu et al.
-2024c) introduces Pairwise-Preference Search (PAIRS) to mitigate biases in LLM evaluators.
-(Wang et al.
-2024b) proposes three methods to calibrate positional bias in LLM-as-a-Judge, including multiple evidence calibration, balanced position calibration, and human-in-the-loop calibration.
-Our evaluation results on scoring bias will serve as a reference to guide future research in mitigating scoring biases in judge models and enhancing the robustness of scoring-based evaluations.
+Liu et al.
+introduces Pairwise-Preference Search (PAIRS) to mitigate biases in LLM evaluators.
+Wang et al.
+proposes three methods to calibrate positional bias in LLM-as-a-Judge, including multiple evidence calibration, balanced position calibration, and human-in-the-loop calibration.
+Our evaluation results on scoring bias will serve as a reference for future research on mitigating scoring biases in judge models and enhancing the robustness of scoring-based evaluations.
 Evaluation Method
-Figure 2: The scoring prompt template with four components from (Kim et al. 2024c). The red text represents the components may be perturbed.
-Figure 1 shows our evaluation framework, inspired by (Ye et al. 2024a).
+Figure 3: Scoring prompt template adapted from (Kim et al. 2024c), with red text indicating components that were perturbed in our experiments.
+Figure 2 shows our evaluation framework, inspired by (Ye et al.
+2024a).
+On a selected dataset, we inject the scoring prompt for the LLM judges with predefined perturbation types.
+Then we obtain the scores from a selected judge model.
+We evaluate the scoring bias by calculating the correlations between judge model’s scores and golden scores, which are obtained from advanced LLMs and human annotators.
+Furthermore, our metrics include score distribution, which is used to measure the impact of different types of bias on the scoring tendencies of LLM judges.
 Dataset Data Size Human Judgment Reference Answer
@@
 Vicuna Bench 320
-Table 1: Statics of our evaluation datasets.
-We evaluate the scoring bias in LLM-as-a-Judge by perturbing the scoring prompt and observing whether the scoring judgments are affected, which follows (Kim et al. 2024c).
-Bias Types
-We define three types of bias targeting the key components related to scoring in LLM judges:
-  • Score Rubrics Order
-  • Score IDs
-  • Reference Answer Score
+Table 1: Statistics of our evaluation datasets.
+Problem Formulation
+We assess the scoring bias in LLM-as-a-Judge by perturbing the scoring prompt and measuring the resulting impact on judgments.
+Following (Kim et al.
+2024c), the input prompt $P$ for the scoring-based LLM judge consists of four components: a task description $T$, a target instance $C^{*}$, which can be either an instruction-response or a question-answer pair, a reference answer with a fixed score $R$, and score rubrics $S$, as shown in Figure 3.
+The input prompt $P$ is formulated as follows:
+$P=(T,C^{*},[R],S),$ (1)
+where * denotes the component will not be perturbed, and $\cdot$] means that the enclosed component is optional within the input prompt.
+This work focuses on biases induced by perturbations on the non-target components of the input prompt for LLM judges.
+Accordingly, we keep the target instance fixed.
+We also treat the reference answer as an optional element to explore its impact on scoring robustness, as it is not a mandatory part of the scoring process.
+By perturbing the reference answer $R\rightarrow R^{\prime}$ or the score rubrics $S\rightarrow S^{\prime}$, we get a perturbed input prompt $P^{\prime}$.
+At the same time, the task description $T$ will also be modified to $T^{\prime}$ in conjunction with the perturbed components to ensure the prompt remains coherent.
+The score from the perturbed prompt will be:
+$s^{\prime}=\textbf{LLM}(P^{\prime}).$ (2)
+For each target instance $C$, we have a golden score $s$ from advanced LLMs or human annotators.
+If $s$ and $s^{\prime}$ differ, it demonstrates that the LLM judge exhibits scoring bias on the given sample.
+Perturbation Types
+We propose three methods to perturb key parts of the judge prompt to introduce scoring biases:
+  • Score Rubric Order: The ordering of the description for each score within the score rubrics. e.g., ascending order (score 1 to score 5), descending order (score 5 to score 1), or random order.
+  • Score ID: Typically, scores are represented using Arabic numerals $\{1,2,3,4,5\}$ as IDs.
+In this work, we explore the use of letter grades $\{\text{E},\text{D},\text{C},\text{B},\text{A}\}$ and Roman numerals $\{\text{i},\text{ii},\text{iii},\text{iv},\text{v}\}$ as scoring IDs.
+  • Reference Answer Score: Based on the finding from (Kim et al.
+2024a, c) that reference answers improve scoring accuracy for LLM judges, our work addresses an unexplored question: whether attaching specific scores to these references could degrade scoring robustness.
 Datasets
@@
 2024b): This is a principled generation benchmark with instance-specific evaluation criteria.
-We use its test set for non-multilingual capabilities with 2,780 responses from 695 instances that are evaluated by human annotators.
+We use the non-multilingual portion of its test set, which contains 2,780 responses for 695 instances, all evaluated by human annotators.
   • FLASK (Ye et al. 2024b): This is a fine-grained evaluation benchmark with 200 test prompts, 12 score rubrics, and 2,001 responses. It also includes scores labeled by human annotators.
@@
 2023): This is a multi-turn chat benchmark with 80 test prompts.
-(Kim et al.
-2024a) write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B (Xu et al.
+Kim et al.
+write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B (Xu et al.
 2024), Vicuna-13B (Chiang et al.
@@
   • Vicuna Bench (Chiang et al.
-2023): This is a single-turn chat benchmark with 80 test prompts.(Kim et al.
-2024a) also write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B, Vicuna-13B, Llama-2-Chat-13B, and GPT-3.5-Turbo-0613.
+2023): This is a single-turn chat benchmark with 80 test prompts.Kim et al.
+also write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B, Vicuna-13B, Llama-2-Chat-13B, and GPT-3.5-Turbo-0613.
 Model BiGGen Bench FLASK
@@
 GPT-4 0.5741 0.6045 0.6207 0.6862
-Table 2: Correlations between LLM judgement pipeline and human when generating golden scores.
-$\rho$ is Spearman’s rank correlation coefficient.
-$r$ is Pearson’s correlation coefficient.
-The GPT-4.1 and GPT-4o scores are generated by our work, the GPT-4 scores are derived from the original datasets.
-The detailed statics of our selected datasets are shown in Table 1.
-For advanced LLM scores, the above datasets have evaluation results obtained from GPT-4 (OpenAI 2023b).
+Table 2: Comparison of the alignment between our LLM judgment pipeline and human golden scores, measured by Spearman’s $\rho$ and Pearson’s $r$ correlation coefficients.
+Scores for GPT-4.1 and GPT-4o are generated using our pipeline, whereas GPT-4 scores are sourced from the original benchmark.
+The detailed statistics of our selected datasets are shown in Table 1.
+Due to the cost of manual annotation, the scores from advanced LLMs are usually used as golden scores (Ye et al.
+2024b; Kim et al.
+2024b).
+The instances in the above four datasets are pre-annotated with scores from GPT-4 (OpenAI 2023b).
 Due to the earlier release of GPT-4, we use more recent models GPT-4.1 (OpenAI 2025) and GPT-4o (OpenAI 2024) to obtain golden scores.
 Inspired by (Kim et al.
-2024a, c), we repeat the scoring three times through a judge prompt with a reference answer that marked a score of 5 and make a majority vote to obtain the final score from LLM evaluators.
-Table 2 shows the correlation between advanced LLM scores and human scores.
-The results show that the GPT-4.1 scores exhibit the highest correlation with human scores, so we select the GPT-4.1 scores as the golden LLM scores.
-Each instance in BiGGen Bench and FLASK is accompanied by a reference answer with score 5.
+2024a, c), we repeat the scoring process three times using a judge prompt that includes a reference answer assigned a score of 5, and then take a majority vote to obtain the final score from the LLM evaluators.
+Table 2 shows that the GPT-4.1 scores exhibit the highest correlation with human scores.
+Throughout the remainder of our experiments, the scores generated by GPT-4.1 are treated as the golden LLM scores.
+Each instance in BiGGen Bench and FLASK is accompanied by a reference answer with a score of 5.
 To measure the impact of the reference answer score on the judgment robustness, it is necessary to additionally generate reference answers with scores ranging from 1 to 4.
@@
 Then a review model checks whether the response matches the corresponding score description.
-If match, we accept the generated responses; otherwise, we rerun the pipeline until a match is achieved.
-We alternately apply GPT-4.1 and GPT-4o as response model and review model in the pipeline.
+If they match, we accept the generated responses; otherwise, we rerun the pipeline until a match is achieved.
+In this pipeline, we use GPT-4.1 and GPT-4o, alternating their roles as the response and review models.
 Metrics
 Figure 4: The proposed generation-review pipeline for synthesizing reference answers.
-To evaluate the scoring bias in LLM-as-a-Judge, we consider the impact or perturbation methods on scoring accuracy and scoring tendency.
+We quantify the effects of scoring bias on both scoring accuracy and scoring tendency.
+Given the golden scores, $S=\{s_{1},\cdots,s_{n}\}$, and the LLM judge’s scores, $S^{\prime}=\{s^{\prime}_{1},\cdots,s^{\prime}_{n}\}$, we quantify scoring accuracy as the correlation between $S$ and $S^{\prime}$.
+Specifically, we report both Spearman’s correlation coefficient $\rho$ (Spearman 1904) and Pearson’s correlation coefficient $r$ (Pearson 1895).
+The scoring tendency is quantified by the score distribution:
+$N_{k}$ $=\sum_{i=1}^{n}{\mathbb{I}(s_{i}=k)},\quad k\in\{1,2,3,4,5\},$ (3)
+$N$ $=(N_{1},N_{2},N_{3},N_{4},N_{5}).$ (4)
+Here, $N_{k}$ represents the count of score $k$ in the scoring results, and $N$ denotes the distribution of the score counts.
 Experiments
-Model Bias Perturbation BiGGen Bench FLASK MT Bench Vicuna Bench
+Model Perturbation Strategy BiGGen Bench FLASK MT Bench Vicuna Bench
 Human GPT-4.1 Human GPT-4.1 GPT-4.1 GPT-4.1
 $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
-GPT-4o - - 0.5685 0.6165 0.7572 0.8059 0.5529 0.5914 0.7369 0.7676 0.7578 0.7724 0.6862 0.7696
+GPT-4o Baseline - 0.5685 0.6165 0.7572 0.8059 0.5529 0.5914 0.7369 0.7676 0.7578 0.7724 0.6862 0.7696
 Rubrics Order Descending 0.5852 0.6245 0.7787 0.8233 0.5754 0.6080 0.7602 0.7760 0.7479 0.7597 0.6734 0.7635
@@
 Roman numerals 0.5754 0.6171 0.7709 0.8148 0.5535 0.5912 0.7380 0.7641 0.7725 0.7786 0.7073 0.7821
-DeepSeek-V3 - - 0.5314 0.5748 0.6775 0.7394 0.5163 0.5660 0.6711 0.7063 0.6323 0.6593 0.4313 0.5443
+DeepSeek-V3 Baseline - 0.5314 0.5748 0.6775 0.7394 0.5163 0.5660 0.6711 0.7063 0.6323 0.6593 0.4313 0.5443
 Rubrics Order Descending 0.5356 0.5812 0.6611 0.7211 0.5181 0.5554 0.6555 0.6858 0.6183 0.6436 0.4993 0.5854
@@
 Roman numerals 0.5164 0.5636 0.6688 0.7195 0.5066 0.5542 0.6750 0.7109 0.6558 0.6642 0.4416 0.5285
-Qwen3-32B - - 0.5421 0.5855 0.7328 0.7715 0.5323 0.5695 0.6549 0.6807 0.6959 0.7053 0.6419 0.7188
+Qwen3-32B Baseline - 0.5421 0.5855 0.7328 0.7715 0.5323 0.5695 0.6549 0.6807 0.6959 0.7053 0.6419 0.7188
 Rubrics Order Descending 0.5572 0.5989 0.7443 0.7862 0.5345 0.5657 0.6586 0.6839 0.6843 0.6989 0.6205 0.7020
@@
 Roman numerals 0.5139 0.5629 0.6992 0.7412 0.5257 0.5781 0.6630 0.6980 0.6360 0.6418 0.6072 0.7126
-Qwen3-8B - - 0.5092 0.5308 0.6522 0.6492 0.5035 0.5309 0.6464 0.6616 0.6238 0.6342 0.6639 0.7284
+Qwen3-8B Baseline - 0.5092 0.5308 0.6522 0.6492 0.5035 0.5309 0.6464 0.6616 0.6238 0.6342 0.6639 0.7284
 Rubrics Order Descending 0.5241 0.5456 0.6458 0.6559 0.4885 0.5339 0.6315 0.6566 0.6005 0.6135 0.5500 0.6647
@@
 Roman numerals 0.4801 0.5120 0.6231 0.6445 0.4606 0.5087 0.6177 0.6468 0.6016 0.5975 0.6111 0.7178
-Mistral-24B - - 0.5059 0.5495 0.6910 0.7201 0.4950 0.5385 0.6303 0.6593 0.5943 0.5762 0.6116 0.6917
+Mistral-24B Baseline - 0.5059 0.5495 0.6910 0.7201 0.4950 0.5385 0.6303 0.6593 0.5943 0.5762 0.6116 0.6917
 Rubrics Order Descending 0.5369 0.5761 0.6987 0.7213 0.4922 0.5327 0.6362 0.6534 0.6104 0.6102 0.6137 0.7082
@@
 Roman numerals 0.5102 0.5590 0.6640 0.6962 0.4957 0.5494 0.6533 0.6872 0.5863 0.5907 0.5903 0.6905
-Table 3: Correlation between golden scores and judge model scores with bias in score rubrics order and score IDs.
-For each judge model, the first row is the metrics without bias, the remaining rows is the metrics under different types of bias.
-The differences between results with and without bias reflect the extent to which bias impacts the scoring robustness.
-DeepSeek-V3 is DeepSeek-V3-671B.
-Mistral-24B is Mistral-Small-24B-Instruct.
-$\rho$ is Spearman’s rank correlation coefficient.
-$r$ is Pearson’s correlation coefficient.
+Table 3: Correlation between golden scores and judge model scores under various prompt perturbations.
+For each judge model, the first row (Baseline) presents the metrics for the original, unperturbed prompt.
+The subsequent rows show the results under different types of perturbations applied to the rubrics order and score IDs.
+The degradation in correlation scores compared to the baseline reflects the model’s scoring bias, indicating its lack of robustness to these perturbations.
+Model names are abbreviated: DeepSeek-V3 denotes DeepSeek-V3-67B and Mistral-24B denotes Mistral-Small-24B-Instruct.
+We report Spearman’s ($\rho$) and Pearson’s ($r$) correlation coefficient.
 Setup
@@
 2024c; Liu et al.
-2024b) has reached a consensus that judge models should be stronger models to ensure evaluation reliability.
-(Ye et al.
-2024a) However, as an exploratory study, we think that capability and size of the judge model should also be considered as influence factors.
+2024b) has reached a consensus that judge models should be stronger models to ensure evaluation reliability (Ye et al.
+2024a).
+After preliminary experiments, we found that scoring bias is a common phenomenon in judge models based on various LLMs.
+Furthermore, we aim to explore the impact of factors such as model type and size on scoring bias.
 Therefore, we select diverse judge models that include closed-source, open-source, large-size, and small-size models.
@@
 2025), and Mistral-Small-24B-Instruct-2501 (MistralAI 2025).
-Those judge models we selected differ from the models used to generate response in the evaluation datasets, thus mitigating potential self-enhancement bias (Zheng et al.
+These judge models we selected differ from the models used to generate responses in the evaluation datasets, thus mitigating potential self-enhancement bias (Zheng et al.
 2023; Ye et al.
@@
 Judge Prompt.
-The prompt we use for scoring-based judgment is shown in Figure 2, following (Kim et al.
-2024c).
+The prompt we use for scoring-based judgment, adapted from (Kim et al.
+2024c), is shown in Figure 3.
 The unperturbed prompt is the one with ascending score rubrics, Arabic numeral score IDs, and no reference answer.
-Hyperparameters.
-As our goals for scoring bias evaluations, we aim to ensure the stability of the model’s generative capabilities. Therefore, we set the judge models’ temperature to 0 while scoring.
+Hyper-parameters. To ensure a deterministic scoring process, which is crucial for a fair evaluation of bias, we set the temperature of the LLM judge to 0 during all scoring tasks.
 Main Results
-Judge models are sensitive to scoring biases in input prompt templates.
-Table 3 and Figure 4 show the correlation of judge model scores and golden scores on all perturbation methods.
-All judge models we evaluated, even for an advanced model such as GPT-4, fluctuations in score results are observed under different biases, indicating that scoring bias has a significant impact on the robustness of LLM-as-a-Judge.
-Furthermore, the impact exhibits certain differences depending on the choice of judge model:
-(1) The capability and size of the judge models are critical factors influencing the robustness against scoring bias.
+Figure 5: Spearman’s correlation between human annotator scores and judge model scores on the FLASK dataset, under perturbations applied to the reference answer scores.
+The existence of scoring bias is proved in all judge models.
+Table 3 and Figure 5 show the correlation between judge model scores and golden scores under all perturbation methods.
+For all judge models we evaluated, including advanced models such as GPT-4o, we observed fluctuations in the scoring results under different types of perturbations.
+This indicates that scoring bias significantly impacts the robustness of LLM-as-a-Judge.
+Furthermore, different judge models exhibit varying sensitivity to scoring bias:
+(1) More powerful judge models tend to be more robust to scoring bias.
 Consistent with existing consensus (Gao et al.
@@
 2024c; Liu et al.
-2024b), our experimental results indicate that judge models with advanced capability and larger size tend to exhibit higher reliability on scoring.
-As shown in Table 3, for example, across all perturbation methods and datasets, the fluctuation in correlation coefficients between GPT-4o and the golden scores remains with in 0.03, except in introduction of a reference answer scored 5.
-By comparing the results of Qwen3-32B and Qwen3-8B, it is evident that judge models with larger parameter sizes demonstrate greater robustness in scoring.
-In comparison with the GPT-4.1 scores in BiGGen-Bench and Vicuna Bench, the correlation coefficients of the Qwen3-8B scores exhibit maximum fluctuations exceeding 0.05 under score rubrics order bias and score IDs bias.
+2024b), our experimental results indicate that judge models with advanced capability and larger size tend to exhibit higher reliability in scoring.
+As shown in Table 3, for example, across all perturbation methods and datasets, the fluctuation in correlation coefficients between GPT-4o and the golden scores remains within 0.03, except for the case involving a reference answer with a score of 5.
+A comparison of the results for Qwen3-32B and Qwen3-8B shows that judge models with larger parameter sizes demonstrate greater robustness in scoring.
+In comparison with the GPT-4.1 scores in BiGGen-Bench and Vicuna Bench, the correlation coefficients of the Qwen3-8B scores exhibit maximum fluctuations exceeding 0.05 under score rubric order bias and score ID bias.
 With reference answer score bias, the maximum fluctuation reaches nearly 0.2.
@@
 Therefore, more capable judge models should be prioritized in scoring evaluations.
-Figure 4: Spearman’s correlation between human annotator scores and judge model scores with bias in reference answer score on FLASK.
-(2) Different judge models exhibit varying sensitivity to scoring bias.
-The impact of bias on the scoring accuracy of judge models can be either positive or negative.
-However, this impact is not fixed; different judge models have different trends in scoring accuracy when affected by the same bias.
-As Table 3 shows, for example, in most cases, the correlation between GPT-4o scores and golden scores increases when Roman numerals are used as scoring IDs.
-But when DeepSeek-V3-671B, Qwen3-32B and Qwen3-8B are employed as judge models, the use of Roman numerals generally exerts a negative effect on their scoring accuracy.
+(2) Contrary to the conventional view that bias is always harmful, we find that scoring bias in LLM-as-a-Judge can also have surprisingly positive effects on the evaluation results.
+For example, as Table 3 shows, in most cases, the correlation between GPT-4o scores and golden scores increases when Roman numerals are used as scoring IDs.
+However, when DeepSeek-V3-671B, Qwen3-32B and Qwen3-8B are employed as judge models, the use of Roman numerals generally exerts a negative effect on their scoring accuracy.
 An interesting phenomenon is that the impact of reference answer score bias on the judge models we evaluated appears to be consistent, which will be further analyzed in subsequent sections.
@@
 (3) Judge models have their own scoring tendencies.
-Figure 5 presents the scoring tendencies of different judge models under various scoring biases on the same dataset FLASK in the form of heatmaps.
+Figure 6 presents the scoring tendencies of different judge models under various scoring biases on the FLASK dataset in the form of heatmaps.
 It is obvious that each judge model exhibits its own specific scoring preferences.
 Under unbiased conditions, DeepSeek-V3-671B assigns a score of 5 to more than half of the instruction-response pairs.
-Compared to the human scoring distribution, Mistral-Small-24B-Instruct has a stronger tendency to score a 4 to the samples.
+Compared to the human scoring distribution, Mistral-Small-24B-Instruct has a stronger tendency to assign a score of 4 to the samples.
 Other judge models show smoother scoring distribution shifts, but certain scores are still preferred.
-GPT-4o demonstrates a preference for score 4, while Qwen3-32B favors score 5.
-An examination of Figure 5 reveals that score rubrics order bias and score IDs bias have a certain impact on the scoring tendencies of the judge models; however, they do not fundamentally alter the overall score distribution.
+GPT-4o demonstrates a preference for a score of 4, while Qwen3-32B favors a score of 5.
+An examination of Figure 6 reveals that score rubric order bias and score ID bias have a certain impact on the scoring tendencies of the judge models; however, they do not fundamentally alter the overall score distribution.
 An exception is Qwen3-8B, whose score distribution is more significantly influenced by biases, possibly due to its small size.
@@
 Therefore, debiasing the judge models is crucial to ensuring the fairness and reliability of the evaluation.
-Figure 5: Heatmaps of judge models’ score distribution with all bias types on FLASK.
-The horizontal axis is the score, where -1 represents the failed judgment.
-The vertical axis is the perturbation method.
-The top lines of all heatmaps are the human scoring tendencies for comparison, and the second lines are results from unbiased judge models.
-Exploratory Experiments and Discussion
-For advanced or large size judge models, application of scoring prompt that differ from usual preferences are worth to be considered.
+Figure 6: Score distributions of LLM judges under various prompt perturbations on the FLASK dataset.
+The horizontal axis represents the assigned score (where -1 indicates a failed judgment), and the vertical axis lists the different perturbation methods applied.
+For reference, the top row in each heatmap (Human) shows the distribution of human scores.
+The second row (None) shows the model’s score distribution for the original, unperturbed prompt.
+Methods for Scoring Bias Mitigation
+For advanced or large size judge models, applying scoring prompt that deviate from conventional preferences is worth considering.
 In almost all scoring-based evaluation tasks, Arabic numerals are commonly used as scoring IDs.
-The score rubrics order typically follows a sequential order from score 1 to score 5.
+The score rubric order typically follows a sequential order from score 1 to score 5.
 These choices align with human writing conventions and cognitive preferences in the design of scoring prompts.
@@
 In more than half of the cases, using letter grades as scoring IDs improves the scoring accuracy of DeepSeek-V3-671B.
-Regarding the score rubrics order bias, adopting a descending order improves the scoring accuracy of GPT-4o, Qwen3-32B, and Mistral-Small-24B-Instruct in most instances.
-Intuitively, employing a random scoring rubrics order often tends to negatively impact the scoring accuracy of all judge models.
-Based on the above analysis, we recommend incorporating certain intentional deviations from conventional practices in the design of scoring prompts.
-The reference answer score bias perform intriguing impacts on the robustness of scoring-based evaluation and play a critical role.
+Regarding the score rubric order bias, adopting a descending order improves the scoring accuracy of GPT-4o, Qwen3-32B, and Mistral-Small-24B-Instruct in most instances.
+Intuitively, employing a random scoring rubric order tends to negatively impact the scoring accuracy of all judge models.
+Based on our analysis, we propose a core design principle for scoring prompts: intentional yet meaningful deviation from convention.
+Effective strategies include introducing structured, unconventional formats, such as presenting scoring rubrics in a descending order or using non-numeric identifiers like letter grades (A, B, C).
+However, the key is to maintain clarity and logical coherence.
+Therefore, chaotic perturbations—such as a fully randomized rubric order or non-hierarchical identifiers—should be avoided.
+Moreover, inspired by previous work (Liu et al.
+2024c; Park et al.
+2024), perturbations applied to scoring prompt also show potential as debiasing strategies.
+For example, effective approaches include implementing multiple variations of the scoring rubric order or IDs and then aggregating the results, or adversarially training the models on samples enhanced with scoring bias perturbations.
+Using full-mark reference answers can stabilize scoring preferences and achieve optimal accuracy.
 The impact of reference answer score bias on the scoring accuracy and scoring tendency of the judge model is presented in Figure 5 and Figure 6.
-The experimental results indicate that, compared to the score rubrics order bias and the score IDs bias, the influence of the reference answer score bias on scoring robustness is both more pronounced and demonstrates stronger regularity.
+The experimental results indicate that, compared to the score rubric order bias and the score ID bias, the influence of the reference answer score bias on scoring robustness is both more pronounced and more regular.
 Figure 5 illustrates the impact of introducing reference answers with different scores in the input prompt on the scoring accuracy of the judge model.
@@
 Specifically, the scoring results from DeepSeek-V3-671B, Qwen3-32B, and Qwen3-8B appear to be highly influenced by the score of the reference answer.
-When the score of the reference answer falls within the range of 1 to 4, these models tend to assign similar scores to the evaluated instruction-response pairs.
+When the score of the reference answer is between 1 and 4, these models tend to assign scores similar to that of the reference answer.
 Interestingly, when a reference answer with a score of 5 is introduced, the scoring behavior of the judge model appears to become more rational.
@@
 However, the introduction of a reference answer scored at 5 mitigates this phenomenon.
-In summary, the above analysis reveals that the score of the reference answer introduced in the prompt has a significant impact on the judge model’s scoring behavior.
+The above analysis reveals that the score of the reference answer introduced in the prompt has a significant impact on the judge model’s scoring behavior.
 In particular, when reference answers with a score of 5 are introduced, the models achieve the highest scoring accuracy without disproportionately influencing their scoring preferences.
-Therefore, it can be concluded that when incorporating reference answers into the scoring-based judge prompt, a full-marked reference answer is the best choice.
+Therefore, it can be concluded that when incorporating reference answers into a scoring-based judge prompt, a full-mark reference answer is the best choice.
 Interestingly, this conceptually aligns with the principles of traditional statistical metrics, such as BLEU and ROUGE, which emphasize the critical role of golden answers in evaluation.
 Limitations and Future Work
-Although our work offers notable conclusions and insights, some limitations still remain for future work.
-Mitigating scoring bias needs further exploration.
-Although there are no dedicated and fancy methods for addressing scoring bias, intuitive and straightforward approaches, such as scoring multiple times followed by majority voting or averaging the scores, require empirical validation for their effectiveness in mitigating scoring bias.
+While our work offers notable conclusions and insights, several limitations remain to be addressed in future work.
+Firstly, our evaluation framework has room for further expansion.
+More types of scoring biases remain to be identified.
+Building more comprehensive benchmarks to evaluate scoring bias in LLM-as-a-Judge is also warranted, as a limited data distribution may lead to an unfair evaluation.
+Secondly, mitigating scoring bias needs further exploration.
+Although there are no dedicated methods for addressing scoring bias, intuitive and straightforward approaches, such as scoring multiple times followed by majority voting or averaging the scores, require empirical validation for their effectiveness in mitigating scoring bias.
+Moreover, the elaborate design of scoring prompt templates, the construction of debiased scoring-based judgment datasets (Park et al.
+2024), and adversarial training are potential approaches to enhancing the scoring robustness of judge models.
 Additionally, although our work demonstrates the existence of scoring bias in LLM-as-a-Judge, the underlying causes of scoring bias remain to be validated.
 Approaches such as training data analysis and information flow observation (Yuan et al.
-2025) may help identify the reason scoring bias exists within or outside the model.
+2025) may help identify the reasons for scoring bias, whether it originates from within the model or from external factors.
 Conclusion
-In this work, we propose scoring bias in LLM-as-a-Judge, defined as scoring judgments shift when judge models are perturbed.
-We design an evaluation framework, including three types of scoring bias: score rubrics order bias, score IDs bias, and reference answer score bias.
-We select four LLM-as-a-Judge benchmarks and expand them as our evaluation datasets through data synthesis.
-Our evaluation metrics can quantify the impact of scoring bias on the scoring accuracy and scoring tendencies of the judge models.
-Our experimental results clearly demonstrate that scoring bias can affect the robustness and reliability of existing scoring-based judge models.
-Judge models with different capabilities, sizes, and architectures exhibit different sensitivities to various scoring biases.
-Our further exploratory experiments and discussions reveal several intriguing phenomena.
-In summary, our work provides detailed evidence and valuable recommendations for evaluating and improving the scoring reliability of LLM-as-a-Judge.
+This work investigates scoring bias in LLM-as-a-Judge, where scoring judgments shift due to minor perturbations in judge prompts.
+To systematically evaluate this, we designed a framework that perturbs score rubric order, score IDs, and reference answers, quantifying the bias using scoring accuracy and scoring tendency.
+Experiments on synthesized datasets expanded from four benchmarks confirm that scoring bias significantly undermines the robustness and reliability of existing LLM judges, with varying sensitivity across different models.
+To mitigate this, we recommend unconventional prompt designs—such as descending rubric order or using letter or Roman numeral IDs—and providing a full-mark reference answer when applicable, both of which effectively improve performance.
+In summary, our work provides systematic evidence and actionable recommendations for evaluating and enhancing the reliability of LLM judges.